lacynowak18:

Arliss has two pieces of carpet runner. One is 2 1/3 yards long and the other is 3 1/3 yards long. She needs 10 yards of carpet runner altogether. How much more does she need to buy?

lacynowak18:

guys sorry for all the question im doing math homework lol

carmelle:

First we do \[2 \frac{ 1 }{ 3 } + 3 \frac{ 1 }{ 3 }\]

carmelle:

What do we get when we do this?

lacynowak18:

idk lol

carmelle:

ok.. lets add the whole numbers

carmelle:

2 + 3 = ?

lacynowak18:

5

carmelle:

ok now for the fractions.. remember to add the numerators only

lacynowak18:

yeah ik i typed it in on my computer and it said it was wrong

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lacynowak18 yeah ik i typed it in on my computer and it said it was wrong \(\color{#0cbb34}{\text{End of Quote}}\) we aren't done yet lol

carmelle:

as I was say, next we do 10 - 5 2/3

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle as I was say, next we do 10 - 5 2/3 \(\color{#0cbb34}{\text{End of Quote}}\) saying*

XxAngelGirlxX:

Solution steps: 1. Add the two given fractions. 2. Subtract the sum from 10 to find your answer.